# Computational Complexity of Sufficiency in Decision Problems

### v2

@@ 3 Computational Complexity of Decision-Relevant Uncertainty
This section establishes the computational complexity of determining which state coordinates are decision-relevant. We prove three main results:
@@
Within the formal model (Section 2) and the succinct encoding used for hardness (Section 2.4), these results sit beyond NP-completeness and provide a rigorous explanation for why over-modeling can be a rational strategy: in the worst case, finding (or even certifying) a minimal set of decision-relevant factors is computationally intractable.
@@ 3.1 Problem Definitions
Problem 3.4 (SUFFICIENCY-CHECK).
@@
Problem 3.5 (MINIMUM-SUFFICIENT-SET).
@@ 3.2 Hardness of SUFFICIENCY-CHECK
SUFFICIENCY-CHECK is coNP-complete [5, 9]. (Machine-verified in Lean 4; see Appendix A.)
@@
SET-COVER MINIMUM-SUFFICIENT-SET Set size maps to coordinate size
@@
Membership in coNP: The complementary problem INSUFFICIENCY is in NP. Given a decision problem $(A,X_{1},\ldots,X_{n},U)$ and coordinate set $I$, a witness for insufficiency is a pair $(s,s^{\prime})$ such that:
@@
    $s_{I}=s^{\prime}_{I}$ (verifiable in polynomial time)
@@
    $\text{Opt}(s)\neq\text{Opt}(s^{\prime})$ (verifiable by evaluating $U$ on all alternatives)
@@ 3.3 Complexity of MINIMUM-SUFFICIENT-SET
MINIMUM-SUFFICIENT-SET is coNP-complete.
@@
Key lemma: A coordinate set $I$ is sufficient if and only if $I$ contains all relevant coordinates (proven formally as sufficient_contains_relevant in Lean):
@@
Consequence: The minimum sufficient set is exactly the set of relevant coordinates. Thus MINIMUM-SUFFICIENT-SET asks: “Is the number of relevant coordinates at most kk?”
@@ 3.4 Anchor Sufficiency (Fixed Coordinates)
Problem 3.8 (ANCHOR-SUFFICIENCY).
@@
ANCHOR-SUFFICIENCY is $\Sigma_{2}^{P}$-complete [14] (already for Boolean coordinate spaces).
@@
If $\exists x^{\star}\,\forall y\,\varphi(x^{\star},y)=1$, then for any $y$ we have $U(\text{YES})=2$ and $U(\text{NO})\leq 1$, so $\text{Opt}(x^{\star},y)=\{\text{YES}\}$ is constant. Conversely, if $\varphi(x,y)$ is false for some $y$, then either $y=0^{m}$ (where NO is optimal) or $y\neq 0^{m}$ (where YES and NO tie), so the optimal set varies across $y$ and the subcube is not constant. Thus an anchor assignment exists iff the $\exists\forall$-SAT instance is true.  $\blacksquare$
@@ 3.6 Implications
Corollary 3.12 (Why Over-Modeling Is Rational).
Finding the minimal set of decision-relevant factors is coNP-complete. Even verifying that a proposed set is sufficient is coNP-complete.
This formally explains the engineering phenomenon:
  1. 1.
    It’s computationally easier to model everything than to find the minimum
  2. 2.
    “Which unknowns matter?” is a hard question, not a lazy one to avoid
  3. 3.
    Bounded scenario analysis (small $\hat{S}$) makes the problem tractable
@@
This connects to the pentalogy’s leverage framework: the modeling budget for deciding what to model is itself a computationally constrained resource.
3.7 Remark: The Collapse to coNP
Early analysis of MINIMUM-SUFFICIENT-SET focused on the apparent $\exists\forall$ quantifier structure, which suggested a $\Sigma_{2}^{P}$-complete result. We initially explored certificate-size lower bounds for the complement, attempting to show MINIMUM-SUFFICIENT-SET was unlikely to be in coNP.
However, the key insight—formalized as sufficient_contains_relevant—is that sufficiency has a simple characterization: a set is sufficient iff it contains all relevant coordinates. This collapses the $\exists\forall$ structure because:
  • The minimum sufficient set is exactly the relevant coordinate set
  • Checking relevance is in coNP (witness: two states differing only at that coordinate with different optimal sets)
  • Counting relevant coordinates is also in coNP
This collapse explains why ANCHOR-SUFFICIENCY retains its $\Sigma_{2}^{P}$-completeness: fixing coordinates and asking for an assignment that works is a genuinely different question. The “for all suffixes” quantifier cannot be collapsed when the anchor assignment is part of the existential choice.
@@ 4 Encoding-Regime Separation
The hardness results of Section 3 apply to worst-case instances under the succinct encoding. This section states an encoding-regime separation: an explicit-state upper bound versus a succinct-encoding worst-case lower bound.
@@
Part 1 is an explicit-state upper bound (time polynomial in $|S|$). Part 2 is a succinct-encoding lower bound under ETH (time exponential in $n$). The encodings are defined in Section 2.4. These two parts are stated in different encodings and are not directly comparable as functions of a single input length.
@@
    Explicit-state upper bound: Under the explicit-state encoding, SUFFICIENCY-CHECK is solvable in time polynomial in $N$ (e.g. $O(N^{2}|A|)$).
@@
    Succinct lower bound under ETH (worst case): Assuming ETH, there exists a family of succinctly encoded instances with $n$ coordinates and minimal sufficient set size $k^{*}=n$ such that SUFFICIENCY-CHECK requires time $2^{\Omega(n)}$.
@@
Part 2 (Succinct ETH lower bound, worst case): A strengthened version of the TAUTOLOGY reduction used in Theorem 3.6 produces a family of instances in which the minimal sufficient set has size $k^{*}=n$: given a Boolean formula $\varphi$ over $n$ variables, we construct a decision problem with $n$ coordinates such that if $\varphi$ is not a tautology then every coordinate is decision-relevant (thus $k^{*}=n$). This strengthening is mechanized in Lean (see Appendix A). Under the Exponential Time Hypothesis (ETH) [8], TAUTOLOGY requires time $2^{\Omega(n)}$ in the succinct encoding, so SUFFICIENCY-CHECK inherits a $2^{\Omega(n)}$ worst-case lower bound via this reduction.  $\blacksquare$

### v3

@@ 4 Computational Complexity of Decision-Relevant Uncertainty
This section establishes the computational complexity of information sufficiency, formalized as coordinate sufficiency, for an arbitrary decision problem $\mathcal{D}$ with factored state space. Because the problems below are parameterized by $\mathcal{D}$, and the $(A,S,U)$ tuple captures any problem with choices under structured information (Remark 1.1), the hardness results are universal: they apply to every problem with coordinate structure, not to a specific problem domain. We prove three main results:
@@
Under the model contract of Section 2.5 and the succinct encoding of Section 2.4, these results place exact relevance identification beyond NP-completeness: in the worst case, finding (or certifying) minimal decision-relevant sets is computationally intractable.
Reading convention for this section.
Each hardness theorem below is paired with a recovery boundary for the same decision question: when structural-access assumptions in Theorem 6.1 hold, polynomial-time exact procedures are recovered; when they fail in [S], the stated hardness applies. (L: CC60, CC36, CC4, CC71.)
@@ 4.1 Problem Definitions
Succinct/circuit encoding.
When $U$ is given succinctly as a Boolean or arithmetic circuit, the instance size refers to the circuit description rather than the (possibly exponentially large) truth table. Evaluating $U(a,s)$ on a single pair $(a,s)$ can be done in time polynomial in the circuit size, but quantification over all states (for example, "for all ss") ranges over an exponentially large domain and is the source of the hardness results below.
@@
Problem 4.4 (SUFFICIENCY-CHECK).
@@
Problem 4.5 (MINIMUM-SUFFICIENT-SET).
@@ 4.2 Hardness of SUFFICIENCY-CHECK
Classical coNP-completeness methodology follows standard NP/coNP reduction frameworks and polynomial-time many-one reducibility [47, 5]. We instantiate that framework for SUFFICIENCY-CHECK with a mechanized TAUTOLOGY reduction.
@@
SUFFICIENCY-CHECK is coNP-complete. (L: CC61, CC60, L18.)
@@
extbfMembership in coNP: The complementary problem INSUFFICIENCY is in NP. A nondeterministic verifier can check a witness pair $(s,s^{\prime})$ as follows:
@@
    Verify $s_{I}=s^{\prime}_{I}$ by projecting and comparing coordinates; this takes time polynomial in the input size.
@@
    Evaluate $U$ on every alternative for both $s$ and $s^{\prime}$. When $U$ is given as a circuit each evaluation runs in time polynomial in the circuit size; using these values the verifier checks whether $\text{Opt}(s)\neq\text{Opt}(s^{\prime})$.
Hence INSUFFICIENCY is in NP and SUFFICIENCY-CHECK is in coNP.
@@
Immediate recovery boundary (SUFFICIENCY-CHECK).
Theorem 4.6 is the [S] general-case hardness statement. For the same sufficiency relation, Theorem 6.1 gives polynomial-time recovery under explicit-state, separable, and tree-structured regimes. (L: CC60, CC71, CC69, CC70, CC72.)
@@ 4.3 Complexity of MINIMUM-SUFFICIENT-SET
MINIMUM-SUFFICIENT-SET is coNP-complete. (L: CC34, CC36.)
@@
Key lemma: In the Boolean-coordinate collapse model, a coordinate set $I$ is sufficient if and only if $I$ contains all relevant coordinates (proven formally as sufficient_iff_relevant_subset / sufficient_iff_relevantFinset_subset in Lean):
@@
Consequence: The minimum sufficient set is exactly the relevant-coordinate set. Thus MINIMUM-SUFFICIENT-SET asks: “Is the number of relevant coordinates at most kk?”
@@
Immediate recovery boundary (MINIMUM-SUFFICIENT-SET).
Theorem 4.7 pairs with Theorem 4.14: exact minimization complexity is governed by relevance-cardinality once the collapse is applied. Recovery then depends on computing relevance efficiently, with structured-access assumptions from Theorem 6.1 giving the corresponding route for the underlying sufficiency computations. (L: CC36, CC34, CC35, S2P2, CC71.)
@@ 4.4 Anchor Sufficiency (Fixed Coordinates)
Problem 4.8 (ANCHOR-SUFFICIENCY).
@@
The second-level hardness template used here follows canonical $\Sigma_{2}^{P}$ polynomial-hierarchy characterizations [47, 5], instantiated to the anchor formulation below.
@@
ANCHOR-SUFFICIENCY is $\Sigma_{2}^{P}$-complete (already for Boolean coordinate spaces). (L: CC5, CC4, L2.)
@@
If $\exists x^{\star}\,\forall y\,\varphi(x^{\star},y)=1$, then for any $y$ we have $U(\text{YES})=2$ and $U(\text{NO})\leq 1$, so $\text{Opt}(x^{\star},y)=\{\text{YES}\}$ is constant. Conversely, fix $x$ and suppose $\exists y_{f}$ with $\varphi(x,y_{f})=0$.
  • If $\varphi(x,0^{m})=1$, then $\text{Opt}(x,0^{m})=\{\text{YES}\}$. The falsifying assignment must satisfy $y_{f}\neq 0^{m}$, where $U(\text{YES})=U(\text{NO})=0$, so $\text{Opt}(x,y_{f})=\{\text{YES},\text{NO}\}$.
  • If $\varphi(x,0^{m})=0$, then $\text{Opt}(x,0^{m})=\{\text{NO}\}$. After padding we have $m>0$, so choose any $y^{\prime}\neq 0^{m}$: either $\varphi(x,y^{\prime})=1$ (then $\text{Opt}(x,y^{\prime})=\{\text{YES}\}$) or $\varphi(x,y^{\prime})=0$ (then $\text{Opt}(x,y^{\prime})=\{\text{YES},\text{NO}\}$). In both subcases the optimal set differs from $\{\text{NO}\}$.
Hence the subcube for this $x$ is not constant. Thus an anchor assignment exists iff the $\exists\forall$-SAT instance is true.  $\blacksquare$
Immediate boundary statement (ANCHOR-SUFFICIENCY).
Theorem 4.9 remains a second-level hardness statement in the anchored formulation; unlike MINIMUM-SUFFICIENT-SET, no general collapse to relevance counting is established here, so the corresponding tractability status remains open in this model. (L: CC5, CC4, L2.)
@@ 4.6 Implications
Corollary 4.12 (Computational Bottleneck for Exact Minimization).
Under the succinct encoding, exact minimization of sufficient coordinate sets is coNP-hard via the $k=0$ case, and fixed-anchor minimization is $\Sigma_{2}^{P}$-complete. (L: CC36, CC4)
Proof.
The $k=0$ case of MINIMUM-SUFFICIENT-SET is SUFFICIENCY-CHECK (Theorem 4.6), giving coNP-hardness for exact minimization. The fixed-anchor variant is exactly Theorem 4.9.  $\blacksquare$
@@
The modeling budget for deciding what to model is therefore a computationally constrained resource under this encoding.[C:Thm. 4.6, Thm. 4.9;R:[S]]
4.7 The succinct–minimal gap and physical agents
In complexity theory, a succinct representation of a combinatorial object is a Boolean circuit of size $\mathrm{poly}(n)$ that computes a function whose explicit (truth-table) form has size $2^{\Theta(n)}$. The term is standard but informal: it denotes a size class (polynomial in input variables), not a structural property. In particular, “succinct” does not mean minimal. The minimal circuit for a function is the smallest circuit that computes it, a rigorous, unique object. A succinct circuit may be far larger than the minimal one while still qualifying as “succinct” by the poly-size criterion.
The gap.
Let $C$ be a succinct circuit computing a utility function $U$, and let $C^{*}$ be the minimal circuit for the same function. The representation gap $\varepsilon=|C|-|C^{*}|$ measures how far the agent’s encoding is from the tightest possible encoding. This gap has no standard formal definition in the literature; it is the unnamed distance between an informal size class and a rigorous structural minimum.
Proposition 4.13 (Coordinate Representation Gap as a Formal Object).
In the Boolean-coordinate model, let $R_{\mathrm{rel}}=\texttt{relevantFinset}(dp)$ and define
$\varepsilon(I)\;:=\;|I\setminus R_{\mathrm{rel}}|+|R_{\mathrm{rel}}\setminus I|.$
Then:
$\varepsilon(I)=0\iff I=R_{\mathrm{rel}}\iff I\text{ is minimal sufficient}.$
(L: S2P3, S2P4, S2P5.)
The results of this paper give $\varepsilon$ formal complexity consequences. By Theorem 4.7, determining the minimal sufficient coordinate set is coNP-complete under the succinct encoding. By Theorem 5.1, under ETH there exist succinct instances requiring $2^{\Omega(n)}$ time to verify sufficiency. In the formal Boolean-coordinate collapse model, size-bounded $\varepsilon=0$ search is exactly the relevance-cardinality predicate (Lean: S2P.min_representationGap_zero_iff_relevant_card). Together: for an agent encoded as a succinct circuit, the worst-case cost of determining whether its own coordinate attention set is minimal is exponential in the number of input coordinates (under ETH). The gap $\varepsilon$ is not merely unnamed; it is computationally irreducible: no polynomial-time algorithm solves all succinct instances (assuming $\mathrm{P}\neq\textsf{coNP}$), and the hard instance family requires $2^{\Omega(n)}$ time.
Physical agents as circuits.
Any physical agent that selects actions based on sensory state can be modeled, in the static snapshot, as a circuit: sensory inputs are coordinates $X_{1},\ldots,X_{n}$; the agent’s internal computation maps states to actions; and the quality of that mapping is measured by a utility function UU.[C:Prop. 2.48, Prop. 2.41;R:[bridge-admissible snapshot]] The explicit-state encoding (truth table) of this mapping has size exponential in the number of sensory dimensions. No physical agent can store it.[C:Thm. 5.1;R:[E] vs [S+ETH]] Every physical agent is therefore a succinct encoding, a circuit whose size is bounded by physical resources (neurons, synapses, transistors, parameters) rather than by the size of the state space it navigates.[C:Thm. 5.1;R:[S]]
This is not a metaphor. A biological nervous system is a feedforward or recurrent circuit that takes sensory state as input and produces motor commands as output.[C:Prop. 2.48;R:[represented adjacent class]] A trained neural network is a parameterized circuit.[C:Prop. 2.48;R:[represented adjacent class]] In both cases, the agent is a succinct encoding of a utility-to-action mapping, and the question “is this agent attending to the right inputs?” is exactly SUFFICIENCY-CHECK under the succinct encoding.[C:Thm. 7.2, Thm. 4.6;R:[S]] The compression that makes a circuit physically realizable, with polynomial size rather than exponential size, is the same compression that makes self-verification exponentially costly in the worst case: the state space the circuit compressed away is precisely the domain that must be exhaustively checked to certify sufficiency.[C:Thm. 5.1, Prop. 5.3;R:[S+ETH],[Q_fin]]
The simplicity tax as $\varepsilon$.
The simplicity tax (Definition 9.2) measures $|\mathrm{Gap}(M,P)|=|R(P)\setminus A(M)|$: the number of decision-relevant coordinates that the agent’s model does not handle natively. In the $\varepsilon$ decomposition above, this is exactly the unattended-relevant component $|R_{\mathrm{rel}}\setminus I|$ (Lean: S2P.representationGap_missing_eq_gapCard, S2P.representationGap_eq_waste_plus_missing). For a physical agent modeled as a circuit, this is the coordinate-level manifestation of $\varepsilon$: the agent attends to a superset of what is necessary, or fails to attend to coordinates that are relevant, and faces worst-case cost exponential in $n$ to verify which case holds (Theorem 4.6). Theorem 9.3 then says the total relevance budget is conserved: unhandled coordinates are not eliminated, only redistributed as external per-site cost.
Consequences for learned circuits.
Large language models and deep networks are succinct circuits with $\mathrm{poly}(n)$ parameters.[C:Thm. 5.1;R:[S]] Mechanistic interpretability asks which internal components (attention heads, neurons, layers) are necessary for a given capability; this is SUFFICIENCY-CHECK applied to the circuit’s internal coordinate structure. Pruning research asks for the minimal subcircuit that preserves performance; this is MINIMUM-SUFFICIENT-SET. The results of this paper imply that these questions are coNP-hard in the worst case under succinct encoding.[C:Thm. 4.6, Thm. 4.7;R:[S]] Empirical progress on interpretability and pruning therefore depends on exploiting the structural properties of specific trained circuits (analogous to the tractable subcases of Theorem 6.1), not on solving the general problem.[C:Thm. 4.6, Thm. 6.1;R:[S] vs [tractable regimes]]
Scope and caveats.
The formal results above apply to the static, deterministic model fixed by C1–C4 (Section 2.5). Biological neural systems add noise, temporal dynamics, and synaptic plasticity; trained networks add stochastic optimization and distribution shift. These factors may alter empirical tractability relative to the static worst-case bounds. The bridge conditions of Proposition 2.40 delineate when static sufficiency results transfer to sequential settings; beyond those conditions, the governing complexity objects change (Propositions 2.432.45). The claims in this subsection are therefore regime-typed consequences of the formal results, not universal assertions about all possible agent architectures.[C:Prop. 2.40, Thm. 2.46, Prop. 2.48;R:[represented adjacent class]]
4.8 Quantifier Collapse for MINIMUM-SUFFICIENT-SET
Theorem 4.14 (Collapse of the Apparent $\exists\forall$ Structure).
The apparent second-level predicate
$\exists I\,(|I|\leq k)\;\forall s,s^{\prime}\in S:\;s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime})$
is equivalent to the coNP predicate $|\text{Relevant}|\leq k$, where
[…]
Consequently, MINIMUM-SUFFICIENT-SET is governed by coNP certificates rather than a genuine $\Sigma_{2}^{P}$ alternation. (L: CC34, CC35, S2P2.)
Proof.
By the formal equivalence sufficient_iff_relevant_subset (finite-set form sufficient_iff_relevantFinset_subset), a coordinate set $I$ is sufficient iff $\text{Relevant}\subseteq I$. Therefore:
$\exists I\;(|I|\leq k\wedge\text{sufficient}(I))\iff\exists I\;(|I|\leq k\wedge\text{Relevant}\subseteq I)\iff|\text{Relevant}|\leq k.$
So the existential-over-universal presentation collapses to counting the relevant coordinates.
A NO certificate for $|\text{Relevant}|\leq k$ is a list of $k+1$ distinct relevant coordinates, each witnessed by two states that differ only on that coordinate and yield different optimal sets; this is polynomially verifiable. Hence the resulting predicate is in coNP, matching Theorem 4.7.
This also clarifies why ANCHOR-SUFFICIENCY remains $\Sigma_{2}^{P}$-complete: once an anchor assignment is existentially chosen, the universal quantifier over the residual subcube does not collapse to a coordinate-counting predicate.  $\blacksquare$
@@ 5 Encoding-Regime Separation
The hardness results of Section 4 apply to worst-case instances under the succinct encoding. This section states an encoding-regime separation: an explicit-state upper bound versus a succinct-encoding worst-case lower bound, and an intermediate query-access obstruction family. The complexity class is a property of the encoding measured against a fixed decision question (Remark 2.21), not a property of the question itself: the same sufficiency question admits polynomial-time algorithms under one encoding and worst-case exponential cost under another.
Physical reading: the separation is an access-and-budget separation for the same decision semantics. Explicit access assumes full boundary observability; succinct access assumes compressed structure that may require exponential expansion to certify universally; query access assumes bounded measurement interfaces.[C:T5.1;P5.3,2.34;R:E,S+ETH,Qf,AR] (L: CC16,CC50,CT6.)
@@
Theorem 5.1 has Part 1 in [E] (time polynomial in $|S|$) and Part 2 in [S+ETH] (time exponential in $n$). Proposition 5.3 is [Q_fin] (finite-state Opt-oracle core), while Propositions 5.7 and 5.11 are [Q_bool] interface refinements (value-entry and state-batch). Relative to Definition 2.23, Propositions 5.8 and 5.15 are explicit simulation-transfer instances. The [E] vs [S+ETH] split in Theorem 5.1 is a same-question encoding/cost-model separation, not a bidirectional simulation claim between those regimes. These regimes are defined in Section 2.4 and are not directly comparable as functions of one numeric input-length parameter.
@@
    [E] Explicit-state upper bound: Under the explicit-state encoding, SUFFICIENCY-CHECK is solvable in time polynomial in $N$ (e.g. $O(N^{2}|A|)$).
@@
    [S+ETH] Succinct lower bound (worst case): Assuming ETH, there exists a family of succinctly encoded instances with $n$ coordinates and minimal sufficient set size $k^{*}=n$ such that SUFFICIENCY-CHECK requires time $2^{\Omega(n)}$.
(L: CC16, CC23, CC24.)
@@
Part 2 (Succinct ETH lower bound, worst case): A strengthened version of the TAUTOLOGY reduction used in Theorem 4.6 produces a family of instances in which the minimal sufficient set has size $k^{*}=n$: given a Boolean formula $\varphi$ over $n$ variables, we construct a decision problem with $n$ coordinates such that if $\varphi$ is not a tautology then every coordinate is decision-relevant (thus $k^{*}=n$). This strengthening is mechanized in Lean (see Appendix A). Under the Exponential Time Hypothesis (ETH) [28, 5], TAUTOLOGY requires time $2^{\Omega(n)}$ in the succinct encoding, so SUFFICIENCY-CHECK inherits a $2^{\Omega(n)}$ worst-case lower bound via this reduction.  $\blacksquare$